| ID | 
Date | 
Author | 
Group | 
Subject | 
Text | 
  |  
| 
  229  
 | 
07 Jun 2023, 15:04 | 
Giuseppe Mercurio, Teguh Asmara | SCS | Morning shift summary | achievements:
 
 
 
	FEL beam realignment because 
pointing was off since we had 0 pulses during | 
   |  | 
  228  
 | 
06 Jun 2023, 23:11 | 
Loic Le Guyader | SCS | late shift summary | achievements
 
 
 
	aligned the beam through 
up to the XOX imager
 
	prepared | 
   |  | 
  227  
 | 
06 Jun 2023, 08:28 | 
Naresh Kujala | XPD | Status | SA2 photon beamline
 
 
- SRA position were 7 mm x 
7 mm, MID reported they see the hallow of | 
     |  | 
  226  
 | 
05 Jun 2023, 08:28 | 
Naresh Kujala | PRC | Status | - SA2 photon beamline
 
 
Today morning found the CRL1: 5 
and 6 arm and Pop-in 45-I are still IN. I | 
    |  | 
  225  
 | 
05 Jun 2023, 06:45 | 
Tommaso Mazza | SQS | Issue | around 6:15, the lights as the one shown 
in the attachment turned on by themselves 
in the control hutch. after some minutes, | 
    |  | 
  224  
 | 
04 Jun 2023, 21:17 | 
Romain Letrun | SPB | Shift summary | X-ray delivery
 
 
 
	12.2-12.45 keV, up to 339 
pulses at 1.13 MHz, 0.8-1.2 mJ.
 
	Couple | 
   |  | 
  223  
 | 
04 Jun 2023, 14:53 | 
Konopkova Zuzana | HED | Shift summary | beam parameters: 20 keV, 300 uJ, 352 pulses
 
 
 
	continuation of user programe.
 
	Machine 
performance very good for trying the 20 keV | 
   |  | 
  222  
 | 
03 Jun 2023, 23:40 | 
Romain Letrun | SPB | Shift summary | X-ray delivery
 
 
 
	10.65 keV, up to 339 pulses 
at 1.13 MHz, 1.6-1.8 mJ.
 
	12.4 | 
   |  | 
  221  
 | 
03 Jun 2023, 13:42 | 
Thomas Preston | HED | Pulse energy has dropped to ~50% from this morning | 03 Jun 2023, 13:46
 
 
We called the BKR to ask what they 
can do about our pulse energy being 50% (~160uJ | 
   |  | 
  220  
 | 
03 Jun 2023, 09:42 | 
Thomas Preston | HED | Shift summary | X-ray parameters: 20 keV,  352 pulses, 
ca 300 uJ
 
 
 
	running user samples.
 
 
 
experiment | 
   |  | 
  219  
 | 
02 Jun 2023, 23:34 | 
Romain Letrun | SPB | Shift summary | X-ray delivery
 
 
 
	10.65 keV, 128 pulses at 
1.13 MHz, 1.6-1.7 mJ.
 
	Overall | 
   |  | 
  218  
 | 
02 Jun 2023, 06:17 | 
Tommaso Mazza |   | pointing stability | we lost spatial overlap on several instances.
 
 
It seems that this happens when 
SASE2 changes number of pulses. | 
    |  | 
  217  
 | 
01 Jun 2023, 23:27 | 
Romain Letrun | SPB | Shift summary | X-ray delivery
 
 
 
	10.65 keV, 128 pulses at 
1.13 MHz and 16 pusles at 141 kHz.
 
	Overall | 
   |  | 
  216  
 | 
01 Jun 2023, 12:10 | 
Konopkova Zuzana | HED | Shift summary | X-ray parameters: 20 keV (20080 eV from 
HIREX), 352 pulses, cca 250 uJ
 
 
 
	changed setup to diamond | 
   |  | 
  215  
 | 
31 May 2023, 23:15 | 
Romain Letrun | SPB | Shift summary | X-ray delivery
 
 
 
	10.65 keV, 128 pulses at 
1.13 MHz and 16 pusles at 141 kHz.
 
	Slope | 
   |  | 
  214  
 | 
31 May 2023, 17:35 | 
Naresh Kujala | PRC | Issue |  
 
 
HED reported that they are unable 
to go 352 pulses and while changing/dailing | 
    |  | 
  213  
 | 
30 May 2023, 23:02 | 
Romain Letrun | SPB | Shift summary | X-ray delivery
 
 
 
	10.7 keV, up to 339 pulses 
at 1.1 MHz.
 
	Large slope through | 
   |  | 
  212  
 | 
28 May 2023, 23:15 | 
Fernando Ardana | FXE | Shift summary | User proposal 3273:
 
 
 
	X-ray delivary
 
	
 
		We 
use several deliveraly modes:
 
		
 
			Continous | 
   |  | 
  211  
 | 
28 May 2023, 22:57 | 
Martin Teichmann | SCS | Shift summary all week | we were trying to make THz and cross correlate 
this to the pump laser.
 
 
We managed to see THz and an auto-correlation | 
   |  | 
  210  
 | 
25 May 2023, 06:35 | 
Tokushi Sato | SPB | Shift summary | X-ray delivery
 
 
 
	10.7 keV, 1.2 mJ, 
1-200 pulses, stable delivery
 
 
 
Optical | 
   |   
 |